OpenStudy (helpisuckatmath):

The absolute value of 4x+5 =2x+4 /4x+5/=2x+4

OpenStudy (phi):

if (we don't know yet) 4x+5 is positive, then the absolute value operation can be dropped (because it does not do anything to a positive number) in other words, one possibility is 4x+5 = 2x+4 we can solve for x: add -2x to both sides 4x-2x + 5 = 2x - 2x + 4 2x+5 = 4 add -5 to both sides 2x = -1 x = -1/2 now check if it works: 4x+5= 4*(-1/2) + 5 = -2+5 = 3 and 2x+4 = 2*(-1/2) + 4 = -1+4 = 3 so it looks good. the other possibility is 4x+5 is negative if we multiply it by -1 we can drop the absolute value operator -4x-5= 2x+4 now solve for x